Prove that the sequence {an/n} is convergent where { an} is a bounded sequence

1
Expert's answer
2022-02-06T13:38:44-0500

Solution:

Given that "\left\{a_{n}\right\}"  is a bounded sequence.

"\Rightarrow \quad\left|a_{n}\right| \leqslant M"  for all "n \in \mathbb{N}" , for some "M \in \mathbb{R}^{+}"

"\Rightarrow\left|\frac{a_{n}}{n}\right| \leqslant \frac{m}{n}"

But the sequence "\frac{m}{n}"  converges to 0 .

"\left.\begin{array}{rl}

\therefore & \lim _{n \rightarrow \infty}\left|\frac{a_{n}}{n}\right| \leqslant \lim _{n \rightarrow \infty} \frac{m}{n}=0 \\

\Rightarrow & \lim _{n \rightarrow \infty}\left|\frac{a_{n}}{n}\right|=0 \quad\left[\because\left|\frac{a_{n}}{n}\right| \text { is a non negative sequence }\right] \\

\Rightarrow & \lim _{n \rightarrow \infty} \frac{a_{n}}{n}=0 \quad\left[\because-\left|\frac{a_{n}}{n}\right| \leqslant \frac{a_{n}}{n} \leqslant\left|\frac{a_{n}}{n}\right|\right. \text{By sandwich theorem}] \\



\end{array}\right]"

Hence, proved.